13.If two events A and B are statistically independent, then: a. As occurrence makes it less likely that B will occur. b. As occurrence makes it more likely that B will occur c. As occurrence makes it impossible that B will occur d. As occurrence has no effect on the probability of B occurring. e. Bs occurrence has minimal effect on the probability of A occurring 14.If two events are independent, then: a. they must be mutually exclusive b. the sum of their probabilities must be equal to one c. the probability of their intersection must be zero d. all of the above e. none of the above 15. Which of the following illustrates the concept of statistically independent? a. P(AIB)=P(A) c. P(Au B) P(A)+P(B) d. a and b only e. all of the above 16. Which of the following pairs of events would seem to be mutually exclusive? a. Vacationing next week in Hawaii and vacationing next week in Paris b. Ordering a hamburger for dinner and ordering French fries for dinner. c. Enrolling in a math class and enrolling in a Spanish class next quarter. d. Majoring in economics and playing on the soccer team. e. Eating a sandwich for lunch and drinking a glass of milk for lunch. 17.Which of the following event pair would seem to be mutually exclusive? a. A studying 12 hours a day. B- sleeping 15 hours a day b. A swimming a mile each day. B-having a full-time job. c. A playing professional basketball. B- making a million dollars. d. A playing Angry Birds. B playing Minecraft. e。As exercising regularly·B-listening to classical music.
Answer #1

Ans.

(13).d. A's occurrence has no effect on the probability of B occurring.

(14).e. None of the above because if two events are independent then P(A and B) =P(A)*P(B)

(15).a. P(A/B) =P(A)

(16).c.

(17).d